OpenStudy (anonymous):

ALSO PLEASE HELP JUST NEED ANSWER QUICK What is the sum of a 17-term arithmetic sequence where the first term is 6 and the last term is -90?

OpenStudy (anonymous):

“You have the right to remain silent. Anything you say can and will be used against you in a court of law. You have the right to an attorney. If you cannot afford an attorney, one will be provided for you. Do you understand the rights I have just read to you? With these rights in mind, do you wish to speak to me?”

OpenStudy (anonymous):

Is it minus 90 or just 90

OpenStudy (anonymous):

its -90 here are the options... -813 -790 -765 -714

OpenStudy (anonymous):

negative 90

OpenStudy (anonymous):

S17= 17/2(6-90) = 17/2(-84)=17(-42)=-714

OpenStudy (anonymous):

thank you sosososo much!! :)